# On the wall-normal velocity variance in canonical wall-bounded turbulence

### v1

@@ 2.3 Scaling parameters
Here, the local shear velocity $u_{\tau z}$ is defined from the total shear stress $\tau(z)$ comprising a turbulent and viscous component. Including the viscous component leads to improved spectral scaling compared to using only the turbulent Reynolds shear stress.
Figure 3: Profiles of the dissipation-based length scale $\ell_{\epsilon}$ defined in Eq. (6): (a) $\ell_{\epsilon}$ in viscous units; (b) $\ell_{\epsilon}$ in outer units; (c) the Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$ representing separation between length scales. The dashed lines in each panel all correspond to $\kappa z$.
@@
An important consideration to the later analysis is scale separation between the production and dissipative regions in the spectrum of turbulence. If $\ell_{\epsilon}$ is the appropriate scale for the production region, this separation can be directly represented by the dissipation-based Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$, where $\eta=(\nu^{3}/\epsilon)^{1/4}$ is the Kolmogorov length scale. From the definitions of $\eta$ and $\ell_{\epsilon}$ in Eq. (6), the Reynolds number can be rewritten as $Re_{\epsilon}=(\ell_{\epsilon}u_{\tau z}/\nu)^{3/4}$. As seen in Fig. 3(c), the value can be approximated as $Re_{\epsilon}\approx(\kappa z^{+})^{3/4}$ in the logarithmic region where $\ell_{\epsilon}\approx\kappa z$ and $u_{\tau z}\sim U_{\tau}$. These findings are consistent with studies that connect the degrees of freedom of a (3-D) turbulence cascade to $(l_{i}/\eta)^{3/4}]^{3}$ (Constantin et al., 1985; Landau & Lifshitz, 2013), when setting the integral scale $L_{i}$ to $l_{\epsilon}$.
@@ 3.1 Spanwise spectrum properties
Example spectra $E_{ww}(k_{y})$ are compared at fixed positions of $z$ in viscous and outer units to evaluate the scaling parameters introduced in §2.3. Then, trends in the peak and low-wavenumber spectrum properties are evaluated across the full boundary layer thickness for all DNS cases. These properties are introduced and visualized in Figure 2(a).
@@
Figure 4 compares the DNS $E_{ww}(k_{y})$ spectra at three positions: $z^{+}=$ 30, $z^{+}=$ 100, and $z/\delta=$ 0.5. The spectra are normalized by the wall-scaling parameters $U_{\tau}$ and $z$ in Fig. 4(a) and by the generalized parameters $u_{\tau z}$ and $\ell_{\epsilon}$ in Fig. 4(b). Normalization with generalized parameters yields a better collapse of the spectrum peak at each of the wall-normal positions shown, indicating that $u_{\tau z}$ and $\ell_{\epsilon}$ are better scales for defining the wavenumber $k_{peak}$ and energy density $E_{peak}$ of the peak across a broad range of Reynolds numbers.
At $z^{+}=$ 30 and $z^{+}=$ 100, the spectral peaks of the highest $Re_{\tau}$ cases are well described by both wall scaling in Fig. 4(a) and the generalized scales in Fig. 4(b). The alignment of both options under high-$Re$ conditions is consistent with the close relation $\ell_{\epsilon}=\kappa z$ seen in Fig. 3. The discrepancies at lower $Re_{\tau}$ in Fig. 4(a) are accounted for using the more general length and velocity scales, indicating that $u_{\tau z}$ and $\ell_{\epsilon}$ can account for finite-Reynolds-number effects in the spectral peak. The only deviation in Fig. 4(b) occurs for the $Re_{\tau}=$ 180 cases at $z^{+}=$ 100, due to the position being in the upper half of the boundary layer where $\delta$-scale effects are substantial.
@@
The peak energy density $E_{peak}$ also follows the generalized scales in Fig. 5(e) more closely than the wall scaling in Fig. 5(b). In particular, $\ell_{\epsilon}$ and $u_{\tau z}$ account for the Reynolds-number effects apparent in Figs. 4(a) and 5(b). The generalized scales also accounts for the higher peak amplitude of the ZPG TBL case seen in Fig. 5(b). The initial increase of $E_{peak}$ within the near-wall viscous region in Fig. 5(e) is likely due in part to insufficient scale separation in the spectrum at these positions, where the dissipative cutoff will directly overlap with and dampen the peak energy at small $Re_{\epsilon}$, i.e. $f(k_{peak}\eta)<1$ in Eq. (3). The normalized energy density decreases moderately in the outer layer before increasing again in the upper half of the boundary layer. The intermediate decrease is not fully understood at this time. The final noteworthy trend in Fig. 5(e) is that the ZPG TBL case has a greater peak in the outer layer, which is related to the low-wavenumber contribution and lower $k_{peak}$ discussed above.
The low-wavenumber limit $E_{k_{y}\to 0}$ in Fig. 5(c,f) follows the same trends as $E_{peak}$ in Fig. 5(b,e). The generalized scales in Fig. 5(e) accounts for the differences in DNS cases within the near-wall region including the logarithmic layer. Consistent with the spectra in Fig. 5(b) at $z/\delta=$ 0.5, the ZPG TBL flow has increased energy relative to $\ell_{\epsilon}$ and $u_{\tau z}$ in the outer layer. The largest $w$ motions are expected to scale with $\delta$ and $U_{\tau}$ based on studies of large-scale-motions in the literature (see, e.g. Guala et al., 2006; Balakumar & Adrian, 2007). However, this expectation is difficult to confirm with the present results because the majority of the $w$ spectrum is governed by the local scale $\ell_{\epsilon}~\sim O(0.1\delta)$ which has limited separation from $\delta$ far from the wall (see, e.g. Fig. 3(b)), and $E_{k_{y}\to 0}$ additionally varies with $z/\delta$ in the outer layer regardless of the normalization.
@@ 3.2 Scale-decomposed variance of the wall-normal velocity
If the $w$ spectrum peak scales with $\ell_{\epsilon}$ and $u_{\tau z}$, these parameters will determine the overall variance corresponding to the integrated spectrum. Two additional factors will contribute to the variance: (1) the largest motions in the given spectrum, which vary in intensity between flow configurations as seen in Fig. 4, and (2) the scale separation between the spectral peak and dissipative cutoff that determines the total small-scale energy. These two factors are evaluated here by computing the area under the spectrum both before and after the peak as classified in Fig. 2(a). The two integrals $\overline{{w^{\prime}}^{2}}_{l}=\int_{0}^{k_{peak}}E_{ww}dk_{y}$ and $\overline{{w^{\prime}}^{2}}_{s}=\int_{k_{peak}}^{\infty}E_{ww}dk_{y}$ sum to the total variance as $\overline{{w^{\prime}}^{2}}=\overline{{w^{\prime}}^{2}}_{l}+\overline{{w^{\prime}}^{2}}_{s}$. In this section, ‘large’ and ‘small’ refer to the size of turbulent motions relative to the spectrum wavenumber peak at a given wall-normal position.
@@
Wall-normal profiles of $\overline{{w^{\prime}}^{2}}_{l}$ and $\overline{{w^{\prime}}^{2}}_{s}$ – based on the integrated spectrum at each $z$ position – are shown in Fig. 6 relative to the local velocity scale $u_{\tau z}$. The vertical axis limits are matched in Fig. 6(a,b) to visualize the relative contribution of the two spectral regions. The contribution of large scales in Fig. 6(a) is relatively small and weakly variant throughout the boundary layer, where the observed range corresponding to the dashed lines is approximately bounded by $\overline{{w^{\prime}}^{2}}_{l}/u_{\tau z}^{2}=$ 0.15 to 0.35. The small range and lack of prominent differences between flow configurations is attributed to two factors: limits in the lowest non-zero wavenumber which results from the domain size, and the method for estimating $\overline{{w^{\prime}}^{2}}_{l}$ which neglects low-wavenumber influences on the spectrum at and beyond the peak $k_{y}\geq k_{peak}$. No further parameterization or interpretation of Fig. 6(a) is given for these reasons.
The high-wavenumber variance contribution $\overline{{w^{\prime}}^{2}}_{s}$ generally aligns with a common curve in the lower half of the boundary layer in Fig. 6(b). The increase with increasing $z$ is expected based on the greater scale separation and more extensive Kolmogorov region of the spectrum farther from the wall. Given the scaling of $k_{peak}$ with $\ell_{\epsilon}$ observed in Fig. 5(d), the Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$ is a more direct representation than $z^{+}$ for the extent of the spectrum beyond the peak. Accordingly, $\overline{{w^{\prime}}^{2}}_{s}$ is plotted versus $Re_{\epsilon}$ in Fig. 6(c), where there is a modest improvement in the collapse of the DNS cases to a common curve. Similar to the limitations for $\overline{{w^{\prime}}^{2}}_{l}$, the estimate of $\overline{{w^{\prime}}^{2}}_{s}$ is imperfect due to the sharp division applied between large- and small-scale contributions.
The local scaling observed in Figs. 6(a,b) can be combined with theoretical forms of the spectrum to propose a semi-empirical expression for the wall-normal velocity variance. At a sufficiently large $Re_{\epsilon}$, an inertial subrange of scales will emerge in the spectrum with self-similar behavior following Kolomogrov’s (1941) hypotheses:
@@
where $C_{w}\approx$ 0.65 is a constant (Saddoughi & Veeravalli, 1994). The lower wavenumber limit of the inertial subrange has been shown to scale with $\ell_{\epsilon}$ (see §2.3 and references therein) such that $E_{ww}=C_{w}\epsilon^{2/3}\ell_{\epsilon}^{5/3}$ at $k_{y}=\ell_{\epsilon}^{-1}$. The observed scaling of the $E_{ww}$ peak with the generalized parameters in Fig. 5, along with the definition in Eq. (6), leads to the same proportional relation:
@@
The definition of $\ell_{\epsilon}$ is therefore inherently consistent with the transition from the larger scales that determine $\epsilon$ to the inertial subrange scales that depend only on $\epsilon$ as seen in Eq. (7). The only external factor in correctly setting this transition is the velocity scale used to define $\ell_{\epsilon}$ in Eq. (6), which must be chosen based on observed scaling in the spectrum.
@@
where $A_{3}=1.5C_{w}$ is relaxed to an undefined constant on account of the neglected peak shape. The start of the integral near $k_{peak}\approx 0.5\ell_{\epsilon}$ also impacts the value of $A_{3}$. The variance within the Kolmogorov region of the spectrum becomes proportional to the velocity scale $u_{\tau z}$ due to its appearance in the definition of $\ell_{\epsilon}$. In addition to setting the length where the -5/3 inertial subrange aligns with the peak scaling, $\ell_{\epsilon}$ also ensures the high-wavenumber variance contribution shares the same velocity scaling as the spectral peak.
Equation (10) is fitted to the DNS cases in Fig. 6(c). The fit includes an additional empirical higher-order term $Re_{\epsilon}^{-2}$ to account for effects at moderate Reynolds numbers when the dissipative cutoff cannot be neglected. Choosing a different (or multiple) higher-order term does not significantly change the goodness of fit. The fit results in the high-wavenumber contribution $\overline{{w^{\prime}}^{2}}_{s}\approx u_{\tau z}^{2}(1.3-1.5C_{w}Re_{\epsilon}^{-2/3}-18Re_{\epsilon}^{-2})$, which aligns well with the DNS up to $z/\delta=$ 0.4 as seen in Fig. 6(c). Combining this fit with the low-wavenumber range $\overline{{w^{\prime}}^{2}}_{l}/u_{\tau z}^{2}=$ 0.15 to 0.35 yields an approximation for the overall variance:
@@
Equation (11) – informed by observed trends in the spectra – indicates that the wall-normal variance should scale with the local shear velocity $u_{\tau z}$ rather than the surface friction velocity $U_{\tau}$. This new scaling is supported by the normalized spectra seen in Figs. 4 and 5. An explanation for the local scaling is given in §4.3.

### v2

@@ 2.3 Flow parameters
Here, the local shear velocity $u_{\tau z}$ is defined from the total shear stress $\tau(z)$ comprising a turbulent and viscous component. Including the viscous component leads to improved agreement with the spectra compared to using only the turbulent Reynolds shear stress.
Figure 3: Profiles of the dissipation-based length $\ell_{\epsilon}$ defined in Eq. (6): (a) $\ell_{\epsilon}$ in viscous units; (b) $\ell_{\epsilon}$ in outer units; (c) the Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$ representing separation between length scales. The dashed lines in each panel all correspond to $\kappa z$.
@@
An important consideration to the later analysis is scale separation between the production and dissipative regions in the spectrum of turbulence. If $\ell_{\epsilon}$ is the appropriate parameter for delineating the production region, this separation can be directly represented by the dissipation-based Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$, where $\eta=(\nu^{3}/\epsilon)^{1/4}$ is the Kolmogorov length scale. From the definitions of $\eta$ and $\ell_{\epsilon}$ in Eq. (6), the Reynolds number can be rewritten as $Re_{\epsilon}=(\ell_{\epsilon}u_{\tau z}/\nu)^{3/4}$. As seen in Fig. 3(c), the value can be approximated as $Re_{\epsilon}\approx(\kappa z^{+})^{3/4}$ in the logarithmic region where $\ell_{\epsilon}\approx\kappa z$ and $u_{\tau z}\sim U_{\tau}$. These findings are consistent with studies that connect the degrees of freedom of a (3-D) turbulence cascade to $(L_{i}/\eta)^{3/4}]^{3}$ (Constantin et al., 1985; Landau and Lifshitz, 2013), when setting the integral scale $L_{i}$ to $\ell_{\epsilon}$.
@@ 3.1 Spanwise spectrum properties
Example spectra $E_{ww}(k_{y})$ are compared at fixed positions of $z$ in viscous and outer units to evaluate whether the parameters introduced in §2.3 can account for the spectrum properties. Then, trends in the peak and low-wavenumber spectrum properties are evaluated across the full boundary layer thickness for all DNS cases. These properties are introduced and visualized in Figure 2(a).
@@
Figure 4 compares the DNS $E_{ww}(k_{y})$ spectra at four positions: $z^{+}=$ 30, $z^{+}=$ 100, $z/\delta=$ 0.1, and $z/\delta=$ 0.5. The spectra are normalized by the wall-scaling parameters $U_{\tau}$ and $z$ in Fig. 4(a) and by the general parameters $u_{\tau z}$ and $\ell_{\epsilon}$ in Fig. 4(b). Normalization with the general parameters yields a better collapse of the spectrum peak at each of the wall-normal positions shown, indicating that $u_{\tau z}$ and $\ell_{\epsilon}$ more accurately define the wavenumber $k_{peak}$ and energy density $E_{peak}$ of the peak across a broad range of Reynolds numbers.
At $z^{+}=$ 30, $z^{+}=$ 100, and $z/\delta=$ 0.1, the spectral peaks of the highest $Re_{\tau}$ cases are well described by both wall scaling in Fig. 4(a) and the general parameters in Fig. 4(b). The alignment of both options under high-$Re$ conditions is consistent with the close relation $\ell_{\epsilon}=\kappa z$ seen in Fig. 3 and previous experimental observations that wall scaling is applicable under these conditions (Kunkel and Marusic, 2006; Baidya et al., 2017). The discrepancies at lower $Re_{\tau}$ in Fig. 4(a) are accounted for using the more general length and velocity parameters, indicating that $u_{\tau z}$ and $\ell_{\epsilon}$ can account for finite-Reynolds-number effects in the spectral peak. Further, the general parameters lead to reduced differences in the spectrum amplitudes of the ZPG and enclosed flows at $z/\delta=$ 0.1. The only deviation in Fig. 4(b) occurs for the $Re_{\tau}=$ 180 cases at $z^{+}=$ 100 and $z/\delta=$ 0.1. At these positions for low Reynolds numbers, both viscous and $\delta$-scale effects are substantial.
@@
The peak energy density $E_{peak}$ also follows the general parameters in Fig. 5(e) more closely than the wall scaling in Fig. 5(b). In particular, $\ell_{\epsilon}$ and $u_{\tau z}$ account for the Reynolds-number effects apparent in Figs. 4(a) and 5(b). The general parameters also account for the higher peak amplitude of the ZPG TBL case seen in Fig. 5(b). The initial increase of $E_{peak}$ within the near-wall viscous region in Fig. 5(e) is likely due in part to insufficient scale separation in the spectrum at these positions, where the dissipative cutoff will directly overlap with and dampen the peak energy at small $Re_{\epsilon}$, i.e. $f(k_{peak}\eta)<1$ in Eq. (3). The normalized energy density decreases moderately in the outer layer before increasing again in the upper half of the boundary layer. The intermediate decrease is not fully understood at this time. The final noteworthy trend in Fig. 5(e) is that the ZPG TBL case has a greater peak in the outer layer, which is related to the low-wavenumber contribution and lower $k_{peak}$ discussed above.
The low-wavenumber limit $E_{k_{y}\to 0}$ in Fig. 5(c,f) follows the same trends as $E_{peak}$ in Fig. 5(b,e). The general parameters in Fig. 5(e) account for the differences in DNS cases within the near-wall region including the logarithmic layer. Consistent with the spectra in Fig. 5(b) at $z/\delta=$ 0.5, the ZPG TBL flow has increased energy relative to $\ell_{\epsilon}$ and $u_{\tau z}$ in the outer layer. The largest $w$ motions are expected to scale with $\delta$ and $U_{\tau}$ based on studies of large-scale-motions in the literature (see, e.g. Guala et al., 2006; Balakumar and Adrian, 2007). However, this expectation is difficult to confirm with the present results because the majority of the $w$ spectrum is governed by the local scale $\ell_{\epsilon}\sim O(0.1\delta)$ which has limited separation from $\delta$ far from the wall (see, e.g. Fig. 3(b)), and $E_{k_{y}\to 0}$ additionally varies with $z/\delta$ in the outer layer regardless of the normalization.
In Fig. 5(d,e,f), there is an abrupt change in the spectral properties at positions far from the wall. The change occurs consistently at $z/\delta\approx$ 0.5 across all cases for each statistic (not shown here). Given that this position corresponds to the peak of $\ell_{\epsilon}$ in Fig. 3(b), it is assumed that $\ell_{\epsilon}$ loses its connection to the spectrum properties as its value decreases in the outer half of the boundary layer.
@@ 3.2 Scale-decomposed variance of the wall-normal velocity
If the $w$ spectrum is proportional to $\ell_{\epsilon}$ and $u_{\tau z}$, these parameters will determine the overall variance corresponding to the integrated spectrum. Two additional factors will contribute to the variance: (1) the largest motions in the given spectrum, which vary in intensity between flow configurations as seen in Fig. 4, and (2) the scale separation between the spectral peak and dissipative cutoff that determines the total small-scale energy. These two factors are evaluated here by computing the area under the spectrum both before and after the peak as classified in Fig. 2(a). The two integrals $\overline{{w^{\prime}}^{2}}_{l}=\int_{0}^{k_{peak}}E_{ww}dk_{y}$ and $\overline{{w^{\prime}}^{2}}_{s}=\int_{k_{peak}}^{\infty}E_{ww}dk_{y}$ sum to the total variance as $\overline{{w^{\prime}}^{2}}=\overline{{w^{\prime}}^{2}}_{l}+\overline{{w^{\prime}}^{2}}_{s}$. In this section, ‘large’ and ‘small’ refer to the size of turbulent motions relative to the spectrum wavenumber peak at a given wall-normal position.
@@
Wall-normal profiles of $\overline{{w^{\prime}}^{2}}_{l}$ and $\overline{{w^{\prime}}^{2}}_{s}$ – based on the integrated spectrum at each $z$ position – are shown in Fig. 6 relative to the local shear velocity $u_{\tau z}$. The vertical axis limits are matched in Fig. 6(a,b) to visualize the relative contribution of the two spectral regions. The contribution of large scales in Fig. 6(a) is relatively small and weakly variant throughout the boundary layer, where the observed range corresponding to the dashed lines is approximately bounded by $\overline{{w^{\prime}}^{2}}_{l}/u_{\tau z}^{2}=$ 0.15 to 0.35. The small range and lack of prominent differences between flow configurations is attributed to two factors: limits in the lowest non-zero wavenumber which results from the domain size, and the method for estimating $\overline{{w^{\prime}}^{2}}_{l}$ which neglects low-wavenumber influences on the spectrum at and beyond the peak $k_{y}\geq k_{peak}$. No further parameterization or interpretation of Fig. 6(a) is given for these reasons.
The high-wavenumber variance contribution $\overline{{w^{\prime}}^{2}}_{s}$ generally aligns with a common curve in the lower half of the boundary layer in Fig. 6(b), and increases sharply in the upper half of the boundary layer where $\ell_{\epsilon}$ loses relevance. The increase with increasing $z$ is expected based on the greater scale separation and more extensive Kolmogorov region of the spectrum farther from the wall. Given the dependence of $k_{peak}$ on $\ell_{\epsilon}$ observed in Fig. 5(d), the Reynolds number $Re_{\epsilon}=\ell_{\epsilon}/\eta$ is a more direct representation than $z^{+}$ for the extent of the spectrum beyond the peak. Accordingly, $\overline{{w^{\prime}}^{2}}_{s}$ is plotted versus $Re_{\epsilon}$ in Fig. 6(c), where there is a modest improvement in the collapse of the DNS cases to a common curve. Similar to the limitations for $\overline{{w^{\prime}}^{2}}_{l}$, the estimate of $\overline{{w^{\prime}}^{2}}_{s}$ is imperfect due to the sharp division applied between large- and small-scale contributions.
The dependence on local parameters observed in Figs. 6(a,b) can be combined with theoretical forms of the spectrum to propose a semi-empirical expression for the wall-normal velocity variance. At a sufficiently large $Re_{\epsilon}$, an inertial subrange of scales will emerge in the spectrum with self-similar behavior following Kolomogrov’s (1941) hypotheses:
@@
where $C_{w}\approx$ 0.65 is a constant (Saddoughi and Veeravalli, 1994). The lower wavenumber limit of the inertial subrange has been shown to be well predicted by $\ell_{\epsilon}$ (see §2.3 and references therein) such that $E_{ww}=C_{w}\epsilon^{2/3}\ell_{\epsilon}^{5/3}$ at $k_{y}=\ell_{\epsilon}^{-1}$. The observed alignment of the $E_{ww}$ peak with the general parameters in Fig. 5, along with the definition in Eq. (6), leads to the same proportional relation:
@@
The definition of $\ell_{\epsilon}$ is therefore inherently consistent with the transition from the larger scales that determine $\epsilon$ to the inertial subrange scales that depend only on $\epsilon$ as seen in Eq. (7). The only external factor in correctly setting this transition is the velocity parameter used to define $\ell_{\epsilon}$ in Eq. (6), which must be chosen based on observed scaling in the spectrum.
@@
where $A_{3}=1.5C_{w}$ is relaxed to an undefined constant on account of the neglected peak shape. The start of the integral near $k_{peak}\approx 0.5\ell_{\epsilon}$ also impacts the value of $A_{3}$. The variance within the Kolmogorov region of the spectrum becomes proportional to the local shear stress and $u_{\tau z}$ due to its appearance in the definition of $\ell_{\epsilon}$. In addition to setting the length where the -5/3 inertial subrange aligns with the peak scaling, $\ell_{\epsilon}$ also ensures the high-wavenumber variance contribution shares the same velocity scaling as the spectral peak.
Equation (10) is fitted to the DNS cases in Fig. 6(c). The fit includes an additional empirical higher-order term $Re_{\epsilon}^{-2}$ to account for effects at moderate Reynolds numbers when the the shape of the dissipative cutoff previously neglected in Eq. (9) has a substantial contribution to the variance. The choice of higher-order exponent is empirical owing to the lack of an analytical solution to the model spectrum and dissipative cutoff used here. Choosing a different (or multiple) higher-order term leads to similar goodness of fit and only small changes to $B_{3}$ that are within the uncertainty range of the low wavenumbers. The fit results in the high-wavenumber contribution $\overline{{w^{\prime}}^{2}}_{s}\approx u_{\tau z}^{2}(1.3-1.5C_{w}Re_{\epsilon}^{-2/3}-18Re_{\epsilon}^{-2})$, which aligns well with the DNS up to $z/\delta=$ 0.4 as seen in Fig. 6(c). Combining this fit with the low-wavenumber range $\overline{{w^{\prime}}^{2}}_{l}/u_{\tau z}^{2}=$ 0.15 to 0.35 yields an approximation for the overall variance:
@@
Equation (11) – informed by observed trends in the spectra – indicates that the wall-normal variance should be computed from the local shear velocity $u_{\tau z}$ rather than the surface friction velocity $U_{\tau}$. This dependence is supported by the normalized spectra seen in Figs. 4 and 5, and an explanation is given in §4.3.